Suppose that in 2015 a country has a population of 1 million and real GDP of $1 billion. In 2016, the population is 1.1 million and the real GDP is $1.1 billion. The real GDP per person growth rate is

1
Expert's answer
2021-11-09T10:45:23-0500

Real GDP per Capita"=\frac{GDP}{Population}"

In 2015: "=\frac{1,000,000,000}{1,000,000}=1,000"

In 2016:"=\frac{1,100,000,000}{1,100,000}=1,000"

The real GDP per person growth rate is:

Percentage change in real GDP per Capita between the two consecutive years:

"=\frac{1,000-1,000}{1,000}\times100"

"=0".